the length of a rectangular plot is 4.5 times that of its breath. if the area of plot is 200 square metres,then what is its length
Answer by TimothyLamb(4379)

---
x = length
y = width
---
x = 4.5y
200 = xy
---
y = 200/x
x = 4.5(200/x)
xx = 900
x = sqrt( 900 )
x = 30
---
answer:
length = 30 m
width = 6.666666 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php